# Taipei 10617, Taiwan

### v2

@@ 2.1 Lie algebra
Each $1d$ Lagrangian ($\ell(\{j_{i}(x^{-})\},x^{-})$ generates an infinitesimal symmetry transformation of the action (2.1)
@@
where $\mathcal{E}$ is a linear operator called the Euler operator, for which $\mathcal{E}(\ell)$ gives the Euler–Lagrange equation for $j(x^{-})$, as derived from the $1d$ Lagrangian ($\ell(\{j_{i}(x^{-})\},x^{-})$, and $\epsilon$ is some infinitesimal parameter. Since the Lagrangian is of finite order, the infinite sum in (2.4) truncates at $k=i$. The Euler operator has the important property that $\mathcal{E}(\ell)=0$ for any value of $j(x^{-})$ and its derivatives if and only if $\ell=\partial_{-}f\sim 0$ is a total derivative olver1993applications . We will use this property repeatedly throughout this paper.
The infinitesimal symmetry transformations (2.4) associated with two Lagrangians $\ell$ and $\ell^{\prime}$ do not in general commute. By direct computation, the commutator $[\delta_{\ell},\delta_{\ell^{\prime}}]\phi$ is
@@
We may write the right hand side of this equation in the form $\frac{1}{2}\mathcal{E}(\ell^{\prime\prime})$, where $\ell^{\prime\prime}$ is a new Lagrangian
@@
where $\ell^{\prime\prime}$ is written in a manifestly anti-symmetric in $(\ell\leftrightarrow\ell^{\prime})$ way, although because Lagrangians are defined modulo total derivatives, we may write ′′=12()()superscript′′$\ell^{\prime\prime}=\frac{1}{2}\partial_{-}\mathcal{E}(\ell)\mathcal{E}(\ell^{\prime})$. The commutator (2.6) defines a Lie algebra on the vector space of $1d$ Lagrangians. If the commutator Lagrangian ′′$\ell^{\prime\prime}=\partial_{-}f\sim 0$ is a total derivative, then $\delta_{\ell^{\prime\prime}}\phi=\frac{1}{2}\mathcal{E}(\ell^{\prime\prime})=0$ is a trivial transformation, and we say that $\ell$ and $\ell^{\prime}$ commute.
@@ 2.2 Korteweg–de Vries type hierarchies
for any $i,j$. Such a subalgebra implies the existence of an infinite hierarchy of integrable models, each model defined by deforming (2.1) with $\ell_{n_{i}}(\{j_{n_{i}}(x^{-})\},x^{-})$
@@
2.3 Mutually commuting subalgebras
There are many mutually commuting subalgebras of (2.6) known so far, of which three are known to be infinite dimensional. We have already mentioned one infinite dimensional subalgebra above, with the first six generators (2.12)–(2.2). Having more explicit examples of mutually commuting subalgebras at our disposal is useful as a guide for understanding their general structure. In this subsection we review the other mutually commuting subalgebras that are known to have at least three generators, listing only the first three generators of each subalgebra for brevity. For a longer list of generators, see Lindwasser:2024qyh .
These subalgebras were found assuming the generators $\ell_{i}$ have definite scaling dimension $[\ell_{i}]$, assigning $x^{-}$ with a scaling dimension $[x^{-}]=-1$ and the current $j(x^{-})$ with some scaling dimension $[j]$, which is not necessarily the canonical scaling dimension 1. In particular, these subalgebras were found with the choices $[j]=0,\pm 1,\pm 2$. The subalgebra associated with the Korteweg–de Vries model (2.12)–(2.2) has definite scaling dimension generators when $[j]=2$.
When $[j]=1$, there are two distinct definite scaling dimension mutually commuting subalgebras (unique up to an affine transformation $j\to aj+b$ with $a,b\in\mathbb{C}$). The first subalgebra corresponds to the symmetries of Liouville, sine-Gordon, and modified Korteweg–de Vries, also proven to be infinite dimensional in Kruskal2
@@ 2.3 Mutually commuting subalgebras
This subalgebra is not known to be infinite dimensional, but is at least six dimensional, with no indication of it terminating at finite dimension. This subalgebra appears to follow the same pattern of orders $n_{i}$ as the subalgebra associated with among other things the Liouville model with generators (2.18)–(2.20). Assuming this subalgebra is infinite dimensional, the lowest order generator $\ell_{1}$ results in the integrable equation
@@
This equation is equivalent after performing potentiation and point transformations to
$\hskip 85.35826ptu_{t}+cu_{x}+d-\frac{3}{2}\frac{u_{x}u_{xx}^{2}}{u_{x}^{2}-1}+u_{xxx}=0$ (2.30)
which belongs to one of the classes of third order integrable equations with constant separant originally found in Svinolupov1982 .
       When $[j]=-1$, there is one distinct definite scaling dimension mutually commuting subalgebra (unique up to an affine transformation $j_{1}\to aj_{1}+b$ with $a,b\in\mathbb{C}$)
@@
This subalgebra is again not known to be infinite dimensional, but is at least six dimensional, with no indication of it terminating at finite dimension. This subalgebra appears to follow the same pattern of orders $n_{i}$ as the subalgebra with generators (2.18)–(2.20) and the subalgebra with generators (2.26)–(2.28). Assuming this subalgebra is infinite dimensional, the lowest order generator $\ell_{1}$ results in the integrable equation
@@
Interestingly, one can show that this equation is also equivalent after performing potentiation and point transformations to (2.30). So although at the level of the Lie algebra, the equations (2.29) and (2.34) are associated with distinct subalgebras, they are of the same class of third order integrable equations in the canonical classification scheme defined up to “almost invertible” transformations SISvinolupov_1992 . This is similar to the relationship between the Korteweg–de Vries and modified Korteweg–de Vries equations, which are related through a Miura transformation.
       And finally when $[j]=-2$, there is one distinct definite scaling dimension mutually commuting subalgebra (unique up to an affine transformation $j_{2}\to aj_{2}+b$ with $a,b\in\mathbb{C}$)
@@
This subalgebra is again not known to be infinite dimensional, but is at least four dimensional, with no indication of it terminating at finite dimension. This subalgebra appears to follow the same pattern of orders $n_{i}$ as the subalgebra associated with the Tzitzéica model with generators (2.22)–(2.3). Assuming this subalgebra is infinite dimensional, the lowest order generator $\ell_{2}$ results in the integrable equation
@@
This equation is equivalent after performing potentiation, point transformations and field redefinitions to
@@
which belongs to one of the classes of fifth order integrable equations with constant separant originally found in Sokolov1985 .
In Lindwasser:2024qyh it was shown that although these subalgebras look quite different from each other, they are in fact related, forming two groups. In particular, the lowest order Lagrangians (2.12), (2.18), (2.26), and (2.31) can all be obtained by taking particular limits of the following order 1 Lagrangian
@@
where $a(j)$ and $b(j)$ are functionals of $j$ such that $a(j)+b(j)^{2}$ is a quartic polynomial in $j$, $d(j)$ is a quadratic polynomial functional of $j$, and $c$ is some constant. This is the most general order 1 Lagrangian which commutes with some order 2 Lagrangian $\ell_{2}$. Meanwhile the lowest order Lagrangians (2.22) and (2.35) can both be obtained by taking particular limits of the following order 2 Lagrangian
@@
where $a$ is a constant, $c(j)$ is a quadratic polynomial functional of $j$, and $b(j,j_{1})$ is a cubic polynomial functional of $j_{1}$
@@
$0\displaystyle 30DD^{\prime\prime}-25D^{\prime 2}+4E^{4}=$ 0 (2.44)
$0\displaystyle 450D^{2}E^{\prime\prime}-150DD^{\prime}E^{\prime}+25D^{\prime 2}E-4E^{5}=$ 0 (2.45)
@@
and $C(j)$ is an undetermined functional. This is the most general order 2 Lagrangian which commutes with some order 3 Lagrangian $\ell_{3}$, but does not commute with an order 1 Lagrangian.
@@ 3 Necessary conditions on Korteweg–de Vries type hierarchies
In this section, we describe some necessary conditions on the generators $\ell_{n_{i}}$ of an infinite dimensional mutually commuting subalgebra of (2.6). We restrict ourselves to consider $1d$ Lagrangians $\ell_{n_{i}}=\ell_{n_{i}}(\{j_{n_{i}}(x^{-})\})$ which commute with the energy-momentum tensor $T=j^{2}$, and hence have no explicit $x^{-}$ dependence. In particular, we will describe the allowed degeneracies of orders of Lagrangians in the subalgebra, as well as the necessary $j_{n_{i}}$ dependence of an order $n_{i}$ Lagrangian $\ell_{n_{i}}$ in such a subalgebra. Some of the following is a summary of results found in Lindwasser:2024qyh .
@@
The equation (3.1) depends on $j(x^{-})$ and its derivatives up to order $2(n_{i}+n_{j})$. Given that we are looking for commuting Lagrangians $\ell_{n_{i}}$ and $\ell_{n_{j}}$ of order $n_{i}$ and $n_{j}$ respectively with $n_{j}\geq n_{i}$ regardless of the value of $j(x^{-})$ and its derivatives, we may regard this equation as a polynomial equation in the variables $j_{n_{j}+1},\dots,j_{2(n_{i}+n_{j})}$. To satisfy this equation for any value of $j(x^{-})$ and its derivatives, the coefficients of each monomial $j_{n_{j}+1}^{N_{1}}\cdots j_{2(n_{i}+n_{j})}^{N_{2n_{i}+n_{j}}}$ must vanish. Recall from Lindwasser:2024qyh that for Lagrangians with $n_{j}>n_{i}>0$, the coefficient of the monomial $j_{2(n_{i}+n_{j})}$ implies
@@
This has the important consequence that in any mutually commuting subalgebra, every generator $\ell_{n_{k}}$ of order $n_{k}>0$ must be quadratic in $j_{n_{k}}$ except for the generators with the lowest nonzero order in the mutually commuting subalgebra.
@@
[…]
@@
comparing the expression in the second line with the commutator (2.6), we see that this constraint means that $\ell_{n_{1}}$ commutes with $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$, where the primes indicate derivatives with respect to $j_{n_{1}}$. This is related to the fact that any integrable equation of the form (1.8) with a potential $V(\{u,u_{x},\dots,u_{2n}\})$ that is order $2n$ has a conserved current $J_{i}$ with time component $J_{t}=\rho_{-1}$ Sokolov1984class
@@
Instead, we have discovered that in any infinite dimensional mutually commuting subalgebra with lowest nonzero order generator $\ell_{n_{1}}$, $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$ commutes with $\ell_{n_{1}}$, and we will see shortly that it is in fact an element of the subalgebra. Because of this, Korteweg–de Vries type hierarchies derived from (1.16) have a conserved current with time component $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$, which after converting to the conventions (1.18) is the same as $\rho_{-1}$. We have shown however that this is a property of any integrable deformation $\delta_{\ell_{n_{i}}}S_{\text{deform}}=0$ of the $2d$ single scalar conformal field theory.
Let us work out what this means with the additional input that $\ell_{n_{1}}$ is the lowest nonzero order generator in the subalgebra. Assuming first that $\ell_{n_{1}}^{\prime\prime\prime}\neq 0$, $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$ is a $1d$ Lagrangian of order $n_{1}$. Recall that when two Lagrangians of the same order commute, they are related via (3.5), and so
@@
where $C$ is some constant, and $f_{n_{1}-1}$ is some order $n_{1}-1$ functional to be determined. Because $\ell_{n_{1}}$ commutes with $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$, (3.10) implies that $\ell_{n_{1}}$ also commutes with $f_{n_{1}-1}$. But we assumed $\ell_{n_{1}}$ has the lowest nonzero order in the mutually commuting subalgebra, and $f_{n_{1}-1}$ has a lower order than $\ell_{n_{1}}$, so $f_{n_{1}-1}=f(j)$ must be an order 0 functional. Without loss of generality, we may remove $f(j)$ from (3.10) by redefining $\ell_{n_{1}}\to\ell_{n_{1}}-f(j)/C$
@@
where now $a_{n_{1}-1}$ and $b_{n_{1}-1}$ are order $n_{1}-1$ functionals and $c\neq 0$ is some constant.
@@
Recall from Lindwasser:2024qyh that the space of order 1 Lagrangians $\ell_{1}$ which commute with an order 2 Lagrangian $\ell_{2}$ takes precisely this form up to some order 0 functional, with the additional constraint that $a(j)+b(j)^{2}$ is a quartic polynomial in $j$.
@@
Recall from Lindwasser:2024qyh that the space of order 2 Lagrangians $\ell_{2}$ which commute with an order 3 Lagrangian $\ell_{3}$ but not an order 1 Lagrangian is up to some order 0 functional
@@
where $b_{1}(j,j_{1})$ is a cubic polynomial in $j_{1}$. This is precisely of the form (3.15), with some additional constraints on the lower order functional $b_{n_{1}-1}$ reviewed in subsection 2.3.
@@
with $f_{n_{1}-1}$ some undetermined order $n_{1}-1$ functional, and $a,b,c$ are constants. Using (3.3), this is precisely the functional dependence of all other generators $\ell_{n_{i}}$ with $n_{i}>n_{1}$ in the subalgebra after replacing $n_{1}\to n_{i}$

### v3

@@ 2.1 Lie algebra
Each $1d$ Lagrangian $\ell(\{j_{i}(x^{-})\},x^{-})$ generates an infinitesimal symmetry transformation of the action (2.1)
@@
where $\mathcal{E}$ is a linear operator called the Euler operator, for which $\mathcal{E}(\ell)$ gives the Euler–Lagrange equation for $j(x^{-})$, as derived from the $1d$ Lagrangian $\ell(\{j_{i}(x^{-})\},x^{-})$, and $\epsilon$ is some infinitesimal parameter. Since the Lagrangian is of finite order, the infinite sum in (2.4) truncates at $k=i$. The Euler operator has the important property that $\mathcal{E}(\ell)=0$ for any value of $j(x^{-})$ and its derivatives if and only if $\ell=\partial_{-}f\sim 0$ is a total derivative olver1993applications . We will use this property repeatedly throughout this paper.
The infinitesimal symmetry transformations (2.4) associated with two Lagrangians $\ell$ and $\ell^{\prime}$ do not, in general, commute. By direct computation, the commutator $\delta_{\ell},\delta_{\ell^{\prime}}]\phi$ is
@@
We may write the right-hand side of this equation in the form $\frac{1}{2}\mathcal{E}(\ell^{\prime\prime})$, where $\ell^{\prime\prime}$ is a new Lagrangian
@@
where $\ell^{\prime\prime}$ is written in a manifestly anti-symmetric in $(\ell\leftrightarrow\ell^{\prime})$ way; although, because Lagrangians are defined modulo total derivatives, we may write $\ell^{\prime\prime}=\frac{1}{2}\partial_{-}\mathcal{E}(\ell)\mathcal{E}(\ell^{\prime})$. The commutator (2.6) defines a Lie algebra on the vector space of $1d$ Lagrangians. If the commutator Lagrangian $\ell^{\prime\prime}=\partial_{-}f\sim 0$ is a total derivative, then $\delta_{\ell^{\prime\prime}}\phi=\frac{1}{2}\mathcal{E}(\ell^{\prime\prime})=0$ is a trivial transformation, and we say that $\ell$ and $\ell^{\prime}$ commute.
@@ 2.2 Korteweg–de Vries type hierarchies
for any $i,j$. Such a subalgebra implies the existence of an infinite hierarchy of integrable models, each model defined by deforming (2.1) with $\ell_{n_{i}}(\{j_{n_{i}}(x^{-})\},x^{-})$,
@@
2.3 Infinite dimensional mutually commuting subalgebras
There are many mutually commuting subalgebras of (2.6) known so far. We have already mentioned one infinite dimensional subalgebra above, with the first six generators (2.12)–(2.2). Having more explicit examples of infinite dimensional mutually commuting subalgebras at our disposal is useful as a guide for understanding their general structure. In this subsection, we review the other infinite dimensional mutually commuting subalgebras, listing only the first three generators of each subalgebra for brevity. In particular, we will confirm in this subsection that the new mutually commuting subalgebras found in Lindwasser:2024qyh are also infinite dimensional.
These subalgebras were found assuming the generators $\ell_{i}$ have definite scaling dimension $\ell_{i}]$, assigning $x^{-}$ with a scaling dimension $x^{-}]=-1$ and the current $j(x^{-})$ with some scaling dimension $[j]$, which is not necessarily the canonical scaling dimension 1. In particular, these subalgebras were found with the choices $j]=0,\pm 1,\pm 2$. The subalgebra associated with the Korteweg–de Vries model (2.12)–(2.2) has definite scaling dimension generators when $[j]=2$.
When $[j]=1$, there are two distinct definite scaling dimension mutually commuting subalgebras (unique up to an affine transformation $j\to aj+b$ with $a,b\in\mathbb{C}$). The first subalgebra corresponds to the symmetries of Liouville, sine-Gordon, and modified Korteweg–de Vries models, also proven to be infinite dimensional in Kruskal2
@@ 2.3 Infinite dimensional mutually commuting subalgebras
This subalgebra appears to follow the same pattern of orders $n_{i}$ as the subalgebra associated with among other things the Liouville model with generators (2.19)–(2.21). The lowest order generator $\ell_{1}$ results in the equation
@@
To show that this equation is integrable, and that it therefore has infinitely many independent symmetries generated by a mutually commuting subalgebra containing the generators (2.27)–(2.29), it is sufficient to show that (2.30) is equivalent to one of the integrable equations in the classification schemes found for instance in Svinolupov1982 . This equation is indeed equivalent after performing potentiation and point transformations to
$\hskip 113.81102ptu_{t}-\frac{3}{2}\frac{u_{x}u_{xx}^{2}}{u_{x}^{2}-1}+u_{xxx}=0$ (2.31)
which belongs to one of the classes of third order integrable equations with constant separant originally found in Svinolupov1982 . We show this in a series of steps, first by making the point transformation $\tilde{u}=1-u^{2}$ on (2.30), resulting in the equation
$\hskip 85.35826pt\tilde{u}_{t}=\partial_{x}\left(-\frac{\tilde{u}_{xx}}{\tilde{u}^{3}}+\frac{3}{4}\frac{3\tilde{u}-2}{(\tilde{u}-1)\tilde{u}^{4}}\tilde{u}_{x}^{2}\right)$ (2.32)
The next step is potentiation, defining $\tilde{u}=\partial_{x}\hat{u}$, we get the equation
$\hskip 85.35826pt\hat{u}_{t}=-\frac{\hat{u}_{xxx}}{\hat{u}_{x}^{3}}+\frac{3}{4}\frac{3\hat{u}_{x}-2}{(\hat{u}_{x}-1)\hat{u}_{x}^{4}}\hat{u}_{xx}^{2}$ (2.33)
Next we make another point transformation, redefining $x=u$ and $\hat{u}=\hat{x}$
$\hskip 85.35826ptu_{t}+\frac{3}{4}\frac{1-2u_{\hat{x}}}{(u_{\hat{x}}-1)u_{\hat{x}}}u_{\hat{x}\hat{x}}^{2}+u_{\hat{x}\hat{x}\hat{x}}=0$ (2.34)
Relabeling $\hat{x}=x$, and performing the last transformation $u\to\frac{1}{2}(u+x)$, we arrive at (2.31).
  When [j]=1[j]=-1, there is one distinct definite scaling dimension mutually commuting subalgebra (unique up to an affine transformation $j_{1}\to aj_{1}+b$ with $a,b\in\mathbb{C}$)
@@
This subalgebra appears to follow the same pattern of orders $n_{i}$ as the subalgebra with generators (2.19)–(2.21) and the subalgebra with generators (2.27)–(2.29). The lowest order generator $\ell_{1}$ results in the equation
@@
Interestingly, one can show that this equation is also equivalent after performing potentiation and point transformations to (2.31). This can be shown in a similar fashion to the steps performed in (2.32)–(2.34). So although at the level of the Lie algebra, the equations (2.30) and (2.38) are associated with distinct subalgebras, they are of the same class of third order integrable equations in the canonical classification scheme defined up to “almost invertible” transformations SISvinolupov_1992 . This demonstrates that (2.38) is integrable, and that it has infinitely many independent symmetries generated by a mutually commuting subalgebra containing the generators (2.35)–(2.37).
  And finally when [j]=2[j]=-2, there is one distinct definite scaling dimension mutually commuting subalgebra (unique up to an affine transformation $j_{2}\to aj_{2}+b$ with $a,b\in\mathbb{C}$)
@@
This subalgebra appears to follow the same pattern of orders $n_{i}$ as the subalgebra associated with the Tzitzéica model with generators (2.23)–(2.3). In this case, the lowest order generator $\ell_{2}$ results in the equation
@@
This equation is integrable, because it is equivalent after performing potentiation, point transformations and field redefinitions to
@@
which belongs to one of the classes of fifth order integrable equations with constant separant originally found in Sokolov1985 . This can again be shown in a similar fashion to the steps performed in (2.32)–(2.34). Because of this, the mutually commuting subalgebra containing the generators (2.39)–(2.41) is infinite dimensional.
In Lindwasser:2024qyh it was shown that, although these subalgebras look quite different from each other, they are in fact related, forming two groups. In particular, the lowest order Lagrangians (2.12), (2.19), (2.27), and (2.35) can all be obtained by taking particular limits of the following order 1 Lagrangian:
@@
where $a(j)$ and $b(j)$ are functionals of $j$ such that $a(j)+b(j)^{2}$ is a quartic polynomial in $j$, $d(j)$ is a quadratic polynomial functional of $j$, and $c$ is some constant. This is the most general order 1 Lagrangian that commutes with some order 2 Lagrangian $\ell_{2}$. Meanwhile the lowest order Lagrangians (2.23) and (2.39) can both be obtained by taking particular limits of the following order 2 Lagrangian
@@
where $a$ is a constant, $c(j)$ is a quadratic polynomial functional of $j$, and $b(j,j_{1})$ is a cubic polynomial functional of $j_{1}$,
@@
$30DD^{\prime\prime}-25D^{\prime 2}+4E^{4}=0$ (2.48)
$450D^{2}E^{\prime\prime}-150DD^{\prime}E^{\prime}+25D^{\prime 2}E-4E^{5}=0$ (2.49)
@@
and $C(j)$ is an undetermined functional. This is the most general order 2 Lagrangian that commutes with some order 3 Lagrangian $\ell_{3}$, but does not commute with an order 1 Lagrangian.
@@ 3 Necessary conditions on Korteweg–de Vries type hierarchies
In this section, we describe some necessary conditions on the generators $\ell_{n_{i}}$ of an infinite dimensional mutually commuting subalgebra of (2.6). We restrict ourselves to consider $1d$ Lagrangians $\ell_{n_{i}}=\ell_{n_{i}}(\{j_{n_{i}}(x^{-})\})$, which commute with the energy-momentum tensor $T=j^{2}$ and hence have no explicit $x^{-}$ dependence. In particular, we will describe the allowed degeneracies of orders of Lagrangians in the subalgebra, as well as the necessary $j_{n_{i}}$ dependence of an order $n_{i}$ Lagrangian $\ell_{n_{i}}$ in such a subalgebra. Some of the following is a summary of results found in Lindwasser:2024qyh .
@@
Equation (3.1) depends on $j(x^{-})$ and its derivatives up to order $2(n_{i}+n_{j})$. Given that we are looking for commuting Lagrangians $\ell_{n_{i}}$ and $\ell_{n_{j}}$ of order $n_{i}$ and $n_{j}$, respectively, with $n_{j}\geq n_{i}$ regardless of the value of $j(x^{-})$ and its derivatives, we may regard this equation as a polynomial equation in the variables $j_{n_{j}+1},\dots,j_{2(n_{i}+n_{j})}$. To satisfy this equation for any value of $j(x^{-})$ and its derivatives, the coefficients of each monomial $j_{n_{j}+1}^{N_{1}}\cdots j_{2(n_{i}+n_{j})}^{N_{2n_{i}+n_{j}}}$ must vanish. Recall from Lindwasser:2024qyh that for Lagrangians with $n_{j}>n_{i}>0$, the coefficient of the monomial $j_{2(n_{i}+n_{j})}$ implies
@@
This has the important consequence that, in any mutually commuting subalgebra, every generator $\ell_{n_{k}}$ of order $n_{k}>0$ must be quadratic in $j_{n_{k}}$ except for the generators with the lowest nonzero order in the mutually commuting subalgebra.
@@
For example, the mutually commuting subalgebras that are so far known to be infinite dimensional are those associated with the Korteweg–de Vries and modified Korteweg–de Vries models, which have generators $\ell_{n_{i}}$ with $n_{1}=1$ for all natural numbers $n_{i}=1,2,3,4,\dots$, i.e. natural numbers satisfying $n_{i}\equiv 0,1\,\text{(mod }2)$, and the Tzitzéica model, which has generators $\ell_{n_{i}}$ with $n_{1}=2$ for all natural numbers $n_{i}$ satisfying $n_{i}\equiv 0,2\,\text{(mod }3)$.
@@
Comparing the expression in the second line with the commutator (2.6), we see that this constraint means that $\ell_{n_{1}}$ commutes with $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$, where the primes indicate derivatives with respect to $j_{n_{1}}$. This is related to the fact that any integrable equation of the form (1.8) with a potential $V(\{u,u_{x},\dots,u_{2n}\})$ that is order $2n$ has a conserved current $J_{i}$ with time component $J_{t}=\rho_{-1}$ Sokolov1984class ,
@@
Instead, we have discovered that in any infinite dimensional mutually commuting subalgebra with lowest nonzero order generator $\ell_{n_{1}}$, $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$ commutes with $\ell_{n_{1}}$, and we will see shortly that it is in fact an element of the subalgebra. Because of this, Korteweg–de Vries type hierarchies derived from (1.19) have a conserved current with time component $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$, which after converting to the conventions (1.21) is the same as $\rho_{-1}$. We have shown, however, that this is a property of any integrable deformation $\delta_{\ell_{n_{i}}}S_{\text{deform}}=0$ of the $2d$ single scalar conformal field theory.
Let us work out what this means with the additional input that $\ell_{n_{1}}$ is the lowest nonzero order generator in the subalgebra. Assuming first that $\ell_{n_{1}}^{\prime\prime\prime}\neq 0$, $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$ is a $1d$ Lagrangian of order $n_{1}$. Recall that, when two Lagrangians of the same order commute, they are related via (3.5), and so
@@
where $C$ is some constant, and $f_{n_{1}-1}$ is some order $n_{1}-1$ functional to be determined. Because $\ell_{n_{1}}$ commutes with $(\ell_{n_{1}}^{\prime\prime})^{-1/(2n_{1}+1)}$, (3.10) implies that $\ell_{n_{1}}$ also commutes with $f_{n_{1}-1}$. But we assumed $\ell_{n_{1}}$ has the lowest nonzero order in the mutually commuting subalgebra, and $f_{n_{1}-1}$ has a lower order than $\ell_{n_{1}}$, so $f_{n_{1}-1}=f(j)$ must be an order 0 functional. Without loss of generality, we may remove $f(j)$ from (3.10) by redefining $\ell_{n_{1}}\to\ell_{n_{1}}-f(j)/C$,
@@
where now $a_{n_{1}-1}$ and $b_{n_{1}-1}$ are order $n_{1}-1$ functionals that come from integration constants when solving (3.11), and $c\neq 0$ is some constant independent of all $j_{k}$, related to $C$.
@@
Recall from Lindwasser:2024qyh that the space of order 1 Lagrangians $\ell_{1}$ that commute with an order 2 Lagrangian $\ell_{2}$ takes precisely this form up to some order 0 functional, with the additional constraint that $a(j)+b(j)^{2}$ is a quartic polynomial in $j$.
@@
Recall from Lindwasser:2024qyh that the space of order 2 Lagrangians $\ell_{2}$ that commute with an order 3 Lagrangian $\ell_{3}$ but not an order 1 Lagrangian is up to some order 0 functional
@@
where $b_{1}(j,j_{1})$ is a cubic polynomial in $j_{1}$. This is precisely of the form (3.15), with some additional constraints on the lower order functional $b_{n_{1}-1}$, reviewed in subsection 2.3.
@@
with $f_{n_{1}-1}$ some undetermined order $n_{1}-1$ functional, and $a,b,c$ are constants. Using (3.3), this is precisely the functional dependence of all other generators $\ell_{n_{i}}$ with $n_{i}>n_{1}$ in the subalgebra after replacing $n_{1}\to n_{i}$,